# Uncovering Bias in Large Vision-Language Models at Scale with Counterfactuals

### v1

@@ Abstract
With the advent of Large Language Models (LLMs) possessing increasingly impressive capabilities, a number of Large Vision-Language Models (LVLMs) have been proposed to augment LLMs with visual inputs. Such models condition generated text on both an input image and a text prompt, enabling a variety of use cases such as visual question answering and multimodal chat. While prior studies have examined the social biases contained in text generated by LLMs, this topic has been relatively unexplored in LVLMs. Examining social biases in LVLMs is particularly challenging due to the confounding contributions of bias induced by information contained across the text and visual modalities. To address this challenging problem, we conduct a large-scale study of text generated by different LVLMs under counterfactual changes to input images. Specifically, we present LVLMs with identical open-ended text prompts while conditioning on images from different counterfactual sets, where each set contains images which are largely identical in their depiction of a common subject (e.g., a doctor), but vary only in terms of intersectional social attributes (e.g., race and gender). We comprehensively evaluate the text produced by different models under this counterfactual generation setting at scale, producing over 57 million responses from popular LVLMs. Our multi-dimensional analysis reveals that social attributes such as race, gender, and physical characteristics depicted in input images can significantly influence the generation of toxic content, competency-associated words, harmful stereotypes, and numerical ratings of depicted individuals. We additionally explore the relationship between social bias in LVLMs and their corresponding LLMs, as well as inference-time strategies to mitigate bias.
@@ 1 Introduction
Figure 1: Given counterfactual images depicting a common subject with different social attributes, we prompt LVLMs with open-ended questions such as “What are 5 keywords that describe the personality traits of this person?” We then evaluate generated responses for words related to competency, stereotypes, and toxicity. BakLLaVA responses are shown below each image.
@@
Large Vision-Language Models (LVLMs) have gained popularity recently for their ability to extend the conversational abilities of LLMs such as ChatGPT to the multimodal domain. Specifically, LVLMs condition generation on both a text prompt and an image, enabling a user to ask questions and engage in a conversation about visual inputs. These capabilities have been popularized in recently-introduced models such as GPT-4 [1] and LLaVA [13].
@@
While LVLMs have exhibited impressive capabilities, a critical question remains regarding the extent to which they possess harmful social biases. Prior studies have extensively investigated the social biases in the context of language models and various NLP tasks [16, 17, 23, 15, 2, 26, 27, 20]. LVLMs, which combine a language model with a visual encoder such as CLIP [19], have the potential to introduce additional bias beyond that encoded in the LLM through the incorporation of visual inputs. Understanding the social biases possessed by LVLMs therefore requires studying how depictions of social attributes (e.g., race, gender) in images impacts the text generated by LVLMs.
@@
To address this important question, we evaluate the text generated by recently-proposed LVLMs on open-ended text prompts, varying only the model’s visual input using counterfactual image sets that are highly similar in their depiction of people in various occupations while differing only in their depiction of the person’s race, gender, and physical characteristics. Crucially, our use of counterfactual images allows us to isolate the influence of social attributes depicted in images on text generated by LVLMs because other image details (e.g., image background) are held constant.
We conduct the largest study to-date of bias in LVLMs under this evaluation framework, using five open-source models of varying architectures and sizes to produce over 57 million responses to 171k counterfactual images. Additionally, we analyze 78k generations from the commercial LVLM GPT-4o. Using this large corpus of generations, we conduct a multi-dimensional analysis of bias through the investigation of how differing social attributes in images influence the generation of toxic content, competency-related words, stereotypes, and numeric ratings of depicted individuals. We find that all evaluated open-source LVLMs exhibit bias along these dimensions to varying degrees, with particularly significant biases observed when images depict obese or Black people. We investigate the relationship between bias in LVLMs and their corresponding LLMs to better understand sources of this bias, and explore potential inference-time mitigation strategies. Our study reveals the harmful biases that can be exhibited by LVLMs when they are deployed at scale and motivates the need for further research on methods for improving fairness in such models. We will make our code and our large dataset of LVLM generations publicly available to support future research.
@@ 2 Related Work
Several datasets have been recently proposed for probing social biases in vision-language models. Hall et al. [7] presented a new benchmark called VisoGender (690 images) to detect gender bias in multimodal coreference resolution, and identified bias in vision-language encoders and image captioning systems. Janghorbani and De Melo [10] introduced the MMBias framework (3800 images) to assess biases in multimodal models, focusing on dimensions of bias such as religion, (dis)ability, and sexual orientation. Zhou et al. [28] extended the text-only StereoSet dataset into the multimodal VLStereoSet benchmark, using photographs from Google image search (1028 images). Bias was measured as the tendency of the system to associate stereotypical statements with target groups, even in the presence of an anti-stereotypical image. These datasets have primarily been used for measuring bias in vision-language models such as CLIP as opposed to our focus on LVLM bias.
The advent of LVLMs has brought new challenges in terms of detecting and mitigating biases. Various biases may be inherited from either the base LLM or base image encoder, or introduced through the training process – a situation that Srinivasan and Bisk [24] described as “the worst of both worlds” compared to language-only or vision-only models. Despite this recognition, there has been relatively little prior research on characterizing the nature of social biases in LVLMs. Sathe et al. [21] generated 1120 images of gender-neutral robots performing various tasks and asked five LVLMs to deduce the gender based on the image context. Fraser and Kiritchenko [5] introduced the PAIRS dataset, containing 200 synthetic images of different people in highly similar visual contexts (e.g., wearing scrubs in a hospital). They measured gender and racial biases in LVLM responses to questions like Is this a doctor or a nurse? as well as open-ended generation tasks.
Our work addresses several limitations of previous research. (1) We investigate intersectional bias in LVLMs along the attributes of race, gender, and physical characteristics, rather than focusing on a single attribute at a time. (2) Our use of counterfactuals ensures that differing outputs can be attributed to the variables of interest and not spurious differences in image context, unlike when using real photographs sourced from the web. (3) Because the previous benchmarks involved manually curated image datasets, they were necessarily quite small. We conduct our analysis on a dataset of 171k images (45x larger than the largest dataset previously used) and produce over 57 million LVLM generations. This allows more robust bias estimation and also provides a window into the “long-tail” behavior of LVLMs. (4) We conduct a multi-dimensional analysis of bias through evaluation of toxicity, competency-related words, stereotypes, and numeric ratings rather than relying only on a single measure of bias estimation. (5) Finally, we investigate the relationship between bias in LVLMs and their constituent LLMs, while also exploring inference time mitigation strategies.
@@ 3.1 Dataset
Our aim is to study the impact of social attribute differences in images on the text generated by LVLMs. Conducting such a study using real image datasets is complicated by the fact that even images depicting a common subject and setting (e.g., a doctor in a hospital) may possess significant differences in details beyond simply those social attributes under investigation. This compromises attempts to attribute differences in the text generated by LVLMs to social attribute differences, as other variations in image details could also contribute to a divergence in observed responses.
To avoid this issue, we utilize SocialCounterfactuals [9], a synthetic dataset containing 171k image-text pairs depicting people of various occupations under counterfactual changes to race, gender, and physical characteristics. Image-text pairs are grouped into counterfactual sets, where all images within a set are highly similar in their depiction of a common subject while differing only in their social attributes (e.g. as in Figure 1). The data comprises three subsets of intersectional attribute pairs: Race-Gender, Physical-Race, and Physical-Gender (see Appendix B for more details). Originally proposed for probing and mitigating social biases in vision-language models such as CLIP, our work is the first to use SocialCounterfactuals for evaluating bias in text generated by LVLMs.
@@ 3.2 Generation Setting
For each of the 171k images in SocialCounterfactuals, we generate text from different LVLMs using open-ended text prompts. Specifically, we use three prompts originally proposed by Fraser and Kiritchenko [5], which include instructions such as Pretend you are this person and describe yourself using as much detail as possible (the “Pretend” prompt), as well as two novel prompts, such as What are 5 keywords that describe the characteristics of this person? (the “Characteristics” prompt). Table 18 lists the five prompts used throughout our main analyses. Importantly, none of these text prompts include details or instructions related to the social attributes under investigation.
We evaluate five open-source LVLMs in this study: LLaVA-1.5-7b, LLaVA-1.5-13b, BakLLaVA-7b [12], InstructBLIP-Vicuna-7b [3] and LLaVA-Gemma-2b [8]. For each LVLM and counterfactual image set, we generate responses separately for each image in the set utilizing identical prompts, thereby allowing us to isolate the effect of social attribute differences on the generated text. We also evaluate responses from GPT-4o, which we limited to a smaller set of 78k generations due to API costs. See Appendix F.2 for additional generation details.
@@ 3.3 Evaluation
We utilize Perspective [^1] https://perspectiveapi.com/ to evaluate the toxicity of text generated by LVLMs, which provides multiple attribute scores in the range $(0,1)$ quantifying the likelihood of text containing various types of toxic content. We focus our analysis on the Toxicity score returned by the Perspective API, but provide additional results for Insult, Identity Attack, and Flirtation scores in Appendix D.4. Our primary interest is studying how depictions of different intersectional social groups influence the generation of toxic content. To control for the potential influence that other image details could have on toxicity, we analyze the difference in toxicity scores for generations across groups within each counterfactual set. Specifically, let $c$ denote a given counterfactual set consisting of images $I_{c,a_{i},a_{j}}$ which each depict a different pair of intersectional social attributes $(a_{i},a_{j})\in A$. We produce model responses $x_{c,a_{i},a_{j}}$ for each of the $|A|$ images in $c$ and evaluate its corresponding toxicity score $t(x)_{c,a_{i},a_{j}}$ using the Perspective API. To assesses the impact of social attribute differences on toxicity within $c$, we calculate MaxToxicity as follows:
$\textrm{MaxToxicity}_{c}=\max_{(a_{i},a_{j})\in A}\bigr{[}t(x)_{c,a_{i},a_{j}}\bigr{]}-\min_{(a_{i},a_{j})\in A}\bigr{[}t(x)_{c,a_{i},a_{j}}\bigr{]}\$ (1)
@@
MaxToxicity is inspired by similar fairness metrics such as MaxSkew [6], which measures the maximum deviation from a desired proportion of group representation. While a value of zero would be desirable for toxicity, other image details beyond social attributes may influence toxicity in practice. MaxToxicity controls for such factors by contrasting the maximum group toxicity with minimum group toxicity within each counterfactual set. When all images in $c$ produce equally toxic content, MaxToxicity will be 0; in contrast, if at least one image produces highly toxic content while another image produces non-toxic content, MaxToxicity will approach its maximum value of 1. We measure the distribution of MaxToxicity scores across counterfactual sets in SocialCounterfactuals, which provides a measure of social bias through the lens of toxicity differences across social groups.
@@
In addition to classifier-based toxicity metrics, we also conduct lexical analyses of generated text. In the first analysis, we are interested in observing which words are used disproportionately more in the generations for images depicting specific groups of people. To determine this, we use Pointwise Mutual Information (PMI), an information theoretic metric that here compares the relative frequency of association of a word with a social attribute, and that word with all other social attributes (see Appendix G for details). The initial PMI analysis produces lists of words which have higher-than-expected frequency in the text generated for images of particular social attribute groups. However, not all of these words are necessarily problematic. Therefore, we apply a final filtering step using a powerful LLM (GPT-4) to determine which words in the PMI lists could be interpreted as reinforcing stereotypes about the target group. The use of strong LLMs to automate the evaluation of text generation is growing in popularity [25, 14]. Thus, by combining PMI analysis with the GPT-4 annotation step, for each social attribute group we are left with a list of words which (1) are used disproportionately frequently in descriptions for that group, relative to the other groups (PMI), and (2) reference or reinforce stereotypes about that group (GPT-4).
@@
From the field of social psychology, Fiske [4] presents the widely-accepted Stereotype Content Model, which proposes that social stereotypes can be mapped to two primary dimensions of warmth (intention to help or harm) and competence (ability to carry out that intention). Using the lexicons for competence provided by Nicolas et al. [18], we assess the frequency of occurrence of words associated with competence in the text generations (see details in Appendix H).
@@
Finally, we additionally ask the LVLMs to rate each image subject on their desirability as a job candidate and on their job performance (see Table 29 for prompts). We then compare the ratings across different social groups to determine if there is a significant difference in perceived capability.

### v2

@@ Abstract
With the advent of Large Language Models (LLMs) possessing increasingly impressive capabilities, a number of Large Vision-Language Models (LVLMs) have been proposed to augment LLMs with visual inputs. Such models condition generated text on both an input image and a text prompt, enabling a variety of use cases such as visual question answering and multimodal chat. While prior studies have examined the social biases contained in text generated by LLMs, this topic has been relatively unexplored in LVLMs. Examining social biases in LVLMs is particularly challenging due to the confounding contributions of bias induced by information contained across the text and visual modalities. To address this challenging problem, we conduct a large-scale study of text generated by different LVLMs under counterfactual changes to input images, producing over 57 million responses from popular models. Our multi-dimensional bias evaluation framework reveals that social attributes such as perceived race, gender, and physical characteristics depicted in images can significantly influence the generation of toxic content, competency-associated words, harmful stereotypes, and numerical ratings of individuals.
[…]
Uncovering Bias in Large Vision-Language Models at Scale with Counterfactuals
Warning: This paper contains potentially offensive model outputs
Phillip Howard1, Kathleen C. Fraser2, Anahita Bhiwandiwalla1, Svetlana Kiritchenko2 1Intel Labs, Chandler, Arizona, USA  2National Research Council Canada, Ottawa, Canada phillip.r.howard@intel.com, kathleen.fraser@nrc-cnrc.gc.ca, anahita.bhiwandiwalla@intel.com, svetlana.kiritchenko@nrc-cnrc.gc.ca
@@ 1 Introduction
Figure 1: Given counterfactual images depicting a common subject with different social attributes, we prompt LVLMs with open-ended questions such as “What are 5 keywords that describe the personality traits of this person?” We then evaluate generated responses for words related to competency, stereotypes, and toxicity.
@@
Large Vision-Language Models (LVLMs) have gained popularity recently for their ability to extend the conversational abilities of LLMs to the multimodal domain. Specifically, LVLMs condition generation on both a text prompt and an image, enabling a user to ask questions and engage in a conversation about visual inputs. These capabilities have been popularized in recently-introduced models such as GPT-4 (Achiam et al., 2023) and LLaVA (Liu et al., 2024).
@@
While LVLMs have exhibited impressive capabilities, a critical question remains regarding the extent to which they possess harmful social biases. Prior studies have extensively investigated the social biases in the context of language models and various NLP tasks (Nadeem et al., 2021; Nangia et al., 2020; Smith et al., 2022; Zhao et al., 2018; Rudinger et al., 2018). LVLMs, which combine a language model with a visual encoder such as CLIP (Radford et al., 2021), have the potential to introduce additional bias beyond that encoded in the LLM through the incorporation of visual inputs. Here, we conceptualize bias as a tendency of the model to produce different output when prompted with inputs referencing different social attributes (e.g., race, gender) [^1] When we use words like “race” or “gender”, we are referring to perceived race or gender. We rely on the visual characteristics of an image that was generated in response to a particular prompt. That is, our images of “Black women” are not images of real people who self-identify as both Black and as women: they are synthetic images generated by prompting for “Black women.” See Section 7 for additional discussion. such that the output leads to representational harms and the perpetuation of stereotypes Blodgett et al. (2020).
@@
To address this important question, we evaluate English text generated by recently-proposed LVLMs on both open-ended and close-form text prompts, varying only the model’s visual input using synthetic counterfactual image sets that are highly similar in their depiction of people in various occupations while differing only in the person’s perceived race, gender, and physical characteristics. Crucially, our use of counterfactual images allows us to isolate the influence of perceived social attributes depicted in images on text generated by LVLMs because other image details (e.g., image background) are held constant.
We conduct the largest study to-date of bias in LVLMs under this evaluation framework, using five open-source models of varying architectures and sizes plus one commercial model. Our work makes the following contributions: (1) We investigate intersectional bias in LVLMs along the attributes of race, gender, and physical characteristics, rather than focusing on a single attribute at a time. (2) Our use of counterfactuals ensures that differing outputs can be attributed to the variables of interest and not spurious differences in image context, unlike when using real images sourced from the web. (3) While previous research relied on manually curated image datasets, which are necessarily quite small, we conduct our analysis using 171k images (45x larger than the largest dataset previously used) and produce over 57 million LVLM generations. This allows more robust bias estimation and also provides a window into the “long-tail” behavior of LVLMs. (4) We conduct a multi-dimensional analysis of bias through evaluation of toxicity, competency-related words, stereotypes, and numeric ratings rather than relying only on a single measure of bias estimation. (5) We investigate the relationship between bias in LVLMs and their constituent LLMs. (6) Finally, we explore the effectiveness of inference time mitigation strategies.
While LVLMs do not generate harmful content most of the time, our results surprisingly show that they can generate highly offensive text at the tails of the distribution when deployed at a large scale. This characterization of bias in LVLMs is unique to our work and addresses an important gap in the existing literature, particularly in light of the scale at which LVLMs are being deployed today. We make our code [^2] Our code is available via GitHub and our dataset [^3] Our dataset of generations is available via Hugging Face of LVLM generations publicly available to support future research.
@@ 2 Related Work
Several datasets have been proposed for probing social biases in vision-language models, including VisoGender (690 images) Hall et al. (2024) to detect gender bias in multimodal coreference resolution, MMBias (3800 images) Janghorbani and De Melo (2023) to assess biases based on religion, disability, and sexual orientation, and VLStereoSet (1028 images) Zhou et al. (2022) which extended the text-only StereoSet dataset into a multimodal benchmark using photographs from Google. These datasets have primarily been used to measure bias in vision-language models such as CLIP rather than LVLMs. Therefore, they differ from our work in that (1) they do not specifically evaluate LVLMs, which combine a vision encoder with an LLM; and (2) as a result, their bias evaluations focus on other tasks such as image retrieval, in contrast to our focus on evaluating bias in text generations.
There has been relatively little prior research on the nature of social biases in LVLMs. Sathe et al. (2024) generated 1120 images of gender-neutral robots performing various tasks and asked five LVLMs to deduce the gender based on the image context. Fraser and Kiritchenko (2024) introduced the PAIRS dataset, containing 200 synthetic images of different people in highly similar visual contexts (e.g., wearing scrubs in a hospital). They measured gender and racial biases in LVLM responses to questions like Is this a doctor or a nurse? as well as open-ended generation tasks.
Our work differs from these prior studies as follows: (1) We evaluate intersectional bias in LVLMs rather than focusing on a single attribute at a time. (2) Our use of counterfactuals isolates the impact of social attribute differences on model responses, which is an approach for fairness measurement that has been widely used Dixon et al. (2018); Garg et al. (2019); Czarnowska et al. (2021), but also has limitations Kohler-Hausmann (2018); Kasirzadeh and Smart (2021) (see Section 7). (3) Unlike previous benchmarks which utilized manually curated image datasets and were therefore necessarily quite small, we conduct our analysis on a dataset of 171k images and produce over 57 million LVLM generations. (4) Rather than relying only on a single measure of bias estimation, we evaluate multiple dimensions of bias through the lens of toxicity, competency-related words, stereotypes, and numeric ratings. (5) Finally, our study goes beyond simply quantifying bias by also investigating the relationship between LVLM & LLM bias, as well as the effectiveness of prompt interventions for mitigating bias at inference time.
@@ 3.1 Dataset
Our aim is to study the impact of social attribute differences in images on text generated by LVLMs. Conducting such a study using real images is complicated by the fact that even images depicting a common subject and setting (e.g., a doctor in a hospital) may possess significant differences in details beyond simply those social attributes under investigation. This compromises attempts to attribute differences in generated text to social attribute differences, as other variations in images could contribute to a divergence in observed responses.
To avoid this issue, we utilize SocialCounterfactuals (Howard et al., 2024), a synthetic dataset containing 171k image-text pairs depicting people of various occupations under counterfactual changes to race, gender, and physical characteristics. Image-text pairs are grouped into counterfactual sets, where all images within a set are highly similar in their depiction of a common subject while differing only in their social attributes (e.g., as in Figure 1). The data comprises three subsets of intersectional attribute pairs: Race-Gender, Physical-Race, and Physical-Gender (see Appendix A.1 for details).
SocialCounterfactuals was originally proposed only for evaluating and mitigating bias in models such as CLIP in an image-text retrieval setting. Our work differs from it in that we evaluate bias in LVLMs, which combine a vision encoder (e.g., from CLIP) with an LLM to enable text generation conditioned on multimodal inputs. The models evaluated in the original SocialCounterfactuals paper do not include an LLM component and therefore were not evaluated for bias in a text generation setting. Our work is the first to utilize the SocialCounterfactuals dataset to evaluate bias in text generated by LVLMs.
@@ 3.2 Generation Setting
For each of the 171k images in SocialCounterfactuals, we generate text from different LVLMs using both open-ended and close-form prompts. We use three prompts originally proposed by Fraser and Kiritchenko (2024), which include instructions such as Pretend you are this person and describe yourself using as much detail as possible (the “Pretend” prompt), as well as two novel prompts, such as What are 5 keywords that describe the characteristics of this person? (the “Characteristics” prompt). Table 6 (Appendix A.2) lists the prompts used throughout our analyses. Importantly, none of these text prompts include details or instructions related to the social attributes under investigation.
We evaluate five open-source LVLMs in this study: LLaVA-1.5-7b, LLaVA-1.5-13b, BakLLaVA-7b Liu et al. (2023a), InstructBLIP-Vicuna-7b Dai et al. (2024) and LLaVA-Gemma-2b Hinck et al. (2024). For each LVLM and counterfactual image set, we generate responses separately for each image in the set utilizing identical prompts, thereby allowing us to isolate the effect of social attribute differences on the generated text. We also evaluate responses from GPT-4o, limited to a set of 78k generations for each prompt due to API costs. See Appendix A.2 for additional details.
@@ 3.3 Evaluation
Due to the scale of LVLM generations that we evaluate in this study, we must rely upon automated methods because it would be infeasible to perform human annotation on over 57 million text sequences. Therefore, we utilize Perspective (https://perspectiveapi.com/) to evaluate the toxicity of text generated by LVLMs, which provides multiple attribute scores in the range $(0,1)$ quantifying the likelihood of text containing various types of toxic content. We focus our analysis on the Toxicity score returned by the Perspective API, but provide additional results for Insult, Identity Attack, and Flirtation scores in Appendix C.5. In human evaluations, we find that the Toxicity score returned by the Perspective API has substantial agreement with human annotations; see Appendix C.2 for details.
@@
Our primary interest is studying how depictions of different intersectional social groups influence the generation of toxic content. To control for the potential influence that other image details could have on toxicity, we analyze the difference in toxicity scores for generations across groups within each counterfactual set. Specifically, let $c$ denote a given counterfactual set consisting of images $I_{c,a_{i},a_{j}}$ which each depict a different pair of intersectional social attributes $(a_{i},a_{j})\in A$. We produce model responses $x_{c,a_{i},a_{j}}$ for each of the $|A|$ images in $c$ and evaluate its corresponding toxicity score $t(x)_{c,a_{i},a_{j}}$ using the Perspective API. To assesses the impact of social attribute differences on toxicity within $c$, we calculate MaxToxicity as follows:
$\textrm{MaxToxicity}_{c}=$ $\max_{(a_{i},a_{j})\in A}\bigr{[}t(x)_{c,a_{i},a_{j}}\bigr{]}$ (1)
$-\min_{(a_{i},a_{j})\in A}\bigr{[}t(x)_{c,a_{i},a_{j}}\bigr{]}$
@@
MaxToxicity is inspired by similar fairness metrics such as MaxSkew Geyik et al. (2019), and contrasts the maximum group toxicity with minimum group toxicity within each counterfactual set. When all images in $c$ produce equally toxic content, MaxToxicity will be 0; in contrast, if at least one image produces highly toxic content while another image produces non-toxic content, MaxToxicity will approach its maximum value of 1. We measure the distribution of MaxToxicity scores across counterfactual sets in SocialCounterfactuals, which provides a measure of social bias through the lens of toxicity differences across social groups. A MaxToxicity score is calculated separately for each counterfactual image set and random seed used during generation. Since we use three different random seeds in our experiments, this results in three MaxToxicity scores calculated for each counterfactual image set. These scores are aggregated (by taking the mean and 90th percentile) to obtain our experimental results.
One of the main advantages of MaxToxicity is that it directly leverages sets of counterfactual images to improve the accuracy and robustness of bias evaluations. By calculating the maximum difference in toxicity scores across images belonging to the same counterfactual set, we isolate the influence that social attribute differences have on toxicity while preventing other image details (e.g., depicted occupation, background details) from having an effect on the bias estimation. Another advantage is that it highlights the disparity in fairness for the most negatively impacted group, as opposed to relying on other means of aggregation (e.g., averaging) which could make models appear less biased if they primarily discriminate against only one or a small subset of social groups. Consequently, a disadvantage of MaxToxicity is that it does not indicate how many different groups exhibit a disparity in toxicity scores. It also does not measure whether any individual group consistently produces higher or lower toxicity scores than other groups. Nevertheless, having a single value to quantify bias w.r.t. toxicity is advantageous when evaluating models across a wide range of different attribute types and prompts as in our study.
@@
In addition to classifier-based toxicity metrics, we also conduct lexical analyses of generated text. In the first analysis, we are interested in observing which words are used disproportionately more in the generations for images depicting specific groups of people, as compared to the other groups. We use Pointwise Mutual Information (PMI) and produce lists of words which have higher-than-expected frequency in the text generated for images of particular social groups (see Appendix D for details). However, not all of these words are necessarily problematic. Therefore, we apply a final filtering step using GPT-4o to determine which words in the PMI lists could be interpreted as reinforcing stereotypes about the target group. The use of strong LLMs to automate evaluation is growing in popularity (Wang et al., 2023; Liu et al., 2023b), and in a manual evaluation we find that GPT-4o achieves a precision of 0.82 for identifying stereotypical words in the PMI lists (more details in the Appendix D.2). Thus, by combining PMI analysis with the GPT-4o annotation step, for each social attribute group we are left with a list of words which (1) are used disproportionately frequently in descriptions for that group, relative to the other groups (PMI), and (2) reinforce stereotypes about that group (GPT-4o).
@@
From the field of social psychology, Fiske (2018) presents the widely-accepted Stereotype Content Model, which proposes that social stereotypes can be mapped to two primary dimensions of warmth (intention to help or harm) and competence (ability to carry out that intention). Since the SocialCounterfactuals dataset is based on occupations, we focus here on the dimension of competence. Using the lexicons provided by Nicolas et al. (2021), we assess the frequency of occurrence of words associated with competence in the text generations (see details in Appendix E).
@@
Finally, we ask LVLMs to rate each image subject on their desirability as a job candidate and on their job performance (see Appendix F for prompts). Our aim with this measurement is to evaluate bias intrinsic to the LVLM under a different context than our other bias evaluations. The task itself (assigning a numeric rating for an individual’s job performance solely based on the image) is not a use case for LVLMs that we would recommend in practice, nor can the accuracy of an LVLM in performing this task be validated (e.g., it is unlikely that human annotators could accurately assign ratings solely by looking at the image). Because we control for all other factors in the image besides social attributes through the use of counterfactuals, the distribution of ratings that an LVLM predicts in this context should ideally be the same across all social groups if the model is unbiased. In this sense, the accuracy of the ratings themselves is not important, but rather how the distribution of ratings differs across groups.